Calculate the market capitalization, earning per share, price/earning ratio, book value per share, and price/book ratio. Thank you

Info:

total stock holder equity: 629

closing share price : 117.51

shares outstanding (in thousands) : 378,597 --> convert to millions since dollars are in millions

below is an income statement for additional information

Net Sales           18,348
Cost of products sold           11,761
Gross Profit             6,587
Marketing, research and general expenses             3,202
Other (income) and expense, net                   27
Operating Profit             3,358
Nonoperating expense                 (59)
Interest income                   10
Interest expense               (318)
Income Before Income Taxes and Equity Interests             2,991
Provision for income taxes               (776)
Income Before Equity Interests             2,215
Share of net income of equity companies                 104
Net Income             2,319
Net income attributable to NCI                 (41)
Net Income Attributable to company A             2,278
Answer #1

MARKET CAPTIALAIZATION=44488.93347 MILLIONS

WORKINGS

MKT. CAPTIALZATION=CURRENT MKT. PER SHARE*TOTAL NO. OF OUTSTANDING SHARES

=117.51*378.597=44488.93347

NO. OF SHARE IN MILLIONS=378597000/1000000=378.597 MILLION

EARNING PER SHARE=6.0170

WORKINGS

EPS=TOTAL EARNING/ NO. OF OUTSTANDING SHARES

=2278/378.597

PRICE/EARNING RATIO=19.5297

WORKING

P/E RATIO=SHARE PRICE/EARNING PER SHARE

=117.51/6.0170=19.5297

BOOK VALE PER SHARE=1.661

WORKING

BOOK VALUE PER SHARE=SHARE HOLDERS EQUITY/NO. OF OUTSTANDING SHARE

=629/378.597=1.661

PRICE/BOOK VALUE=70.7465

WORKINGS

P/B VALUE=MARKET PRICE SHARE/BOOK VALUE PER SHARE

=117.51/1.661=70.7465